OpenStudy (anonymous):

Find an equation of a hyperbola with a=452 units and c=765 units. Assume that the transverse axis is horizontal. a.x^2/204,304 *y^2/380,921=1 b.x^2/380,921*y^2/204,304=1 c.x^2/452*y^2/313=1 d.x^2/313-y^2/452=1

OpenStudy (anonymous):

well this is as simple as pythagorean thm a^2 +b^2 = c^2 --> b^2 = 765^2 - 452^2 hyperbola equation : x2a2−y2b2=1 *dumbcow

OpenStudy (anonymous):

\[b^2=c^2-a^2=380921 \]

OpenStudy (anonymous):

It should be a.

OpenStudy (anonymous):

\[ \frac {x^2}{a^2}- \frac {y^2}{b^2}= 1\\ \frac {x^2}{204304}- \frac {y^2}{380921}= 1\\ \]